a box contains 9 plain pencils and 3 pens. A second box contains 3 color pencils and 5 crayons. One item from each box is chosen at random. What is the probability that a pen from the first box and a crayon from the second box are selected?
2 You have 5 chances in 8 of pickin a colored pencil ===> 5/8
3. The probability of doing both is the product of the two ===> 1/4 * 5/8 = 5/32